### v1

@@ 1 Introduction
Empirical evaluation is the primary mechanism through which progress in foundation models is recognized, compared, and acted upon. As machine learning has shifted from narrow, task-specific systems to general-purpose vision-language models (VLMs) with broad and compositional capabilities (wei2022emergentabilitieslargelanguage), benchmarks now play an outsized role: they define what counts as progress and directly shape how substantial computational and human resources are allocated. Evaluations are no longer a passive reporting tool but an active driver of research direction.
@@
However, modern evaluation pipelines are increasingly misaligned with the behaviors they aim to measure. As model inputs span multiple modalities and outputs become increasingly generative and stochastic, benchmarks must better disentangle genuine capabilities from superficial heuristics and inherent variance (stochastic_variation_lu2022fantasticallyorderedpromptsthem). While the evaluation of language-only models has received sustained methodological attention (llm_eval_mature_srivastava2023imitationgamequantifyingextrapolating; llm_eval_mature_2023opencompass), VLM evaluation remains comparatively under-examined.
@@
Recent evidence suggests that this gap has become a serious liability. Existing VLM benchmarks suffer from pervasive data quality failures, including mislabeled or ambiguous examples, questions solvable without visual input, and heavy reliance on multiple-choice formats that are not representative of downstream use-cases and are vulnerable to spurious correlations (chartqapro; liu2024mmbench; realworldqa2024; zhang2024mme; yang2024ccocrcomprehensivechallengingocr). These artifacts inflate reported accuracy, introduce a substantial noise floor, and reduce the signal-to-noise ratio of the evaluations. In such a regime, small improvements, often on the order of a few percent, are more plausibly explained by overfitting to benchmark idiosyncrasies than by real capability gains, rendering the research community vulnerable to hill-climbing on noise (hillclimb_noise_recht2019imagenetclassifiersgeneralizeimagenet; hillclimb_noise_wei2023inversescalingushaped).
@@
At the same time, evaluation has become a major computational bottleneck. Running comprehensive VLM evaluation suites now consumes a nontrivial fraction of total development compute (strubell2019energypolicyconsiderationsdeep). For example, during the development of OLMo3, nearly 20% of the total compute budget for the post-training phase was reportedly dedicated to evaluation alone (lambert2025goodresearchers). This burden is amplified for VLMs by the dense visual token sequences required to represent high-resolution images and the extended reasoning traces at inference time, which can collectively exceed tens of thousands of tokens per example (qwen3vl). Detailed analyses indicate that much of this cost is spent evaluating samples that are either trivial, noisy, or weakly discriminative (fluid2025; polo2024tinybenchmarks).
@@
In this work, we argue that the design of effective evaluation should be treated as a data curation problem. Rather than repeatedly constructing new benchmarks from scratch, we propose to systematically transform and filter evaluation data to maximize faithfulness, discriminative power, and efficiency. This perspective mirrors recent successes in training data curation, in which careful data transformation and selection has produced large gains in model quality and compute efficiency (fang2023datafilteringnetworks; pmlr-v202-joshi23b; abbas2023semdedupdataefficientlearningwebscale; pmlr-v238-joshi24a; joshi2025datasetdistillationknowledgedistillation; joshi2025mmgenenhancingtaskperformance; datologyai_clip_2024; datologyai_text_2024; beyondweb). We show that the same principles apply, with similar impact, to evaluation.
@@
First, multiple-choice formats are both unfaithful and weakly discriminative in generative settings. Converting MCQs to open-ended generation reveals large hidden capability gaps. On AI2D, for example, average accuracy drops from 77.56% to 40.53%, with the strongest MCQ model losing nearly 35 points. When generative conversion is infeasible, circular evaluation (liu2024mmbench) collapses chance baselines and exposes similar inflation effects.
Second, many VLM benchmarks can be solved without vision. By evaluating models with the image removed, we find that over 70% of samples in VQA-v2 (balanced_vqa_v2) can be answered correctly using language priors alone. Such examples fundamentally fail to measure multimodal reasoning.
Third, low-resolution inputs and inaccurate or ambiguous annotations introduce substantial noise. Using a multi-stage filtering pipeline, we discard up to 42.07% of samples in benchmarks such as MME-RealWorld (Autonomous Driving) (zhang2024mme). In these instances, evaluation is confounded by factual labeling errors and indeterminable ground truths—where poor image quality renders the target objects unrecognizable even to a human observer—effectively precluding reliable performance assessment.
@@ 2 Related Work
Faithful Evaluation. Recent research has identified significant issues with the validity of VLM benchmarks, prompting various mitigation strategies. To address inflated performance caused by high-risk baselines in multiple-choice evaluations, several studies propose reformulating tasks into generative answer-matching settings (chandak2025answer) or employing circular evaluation techniques (liu2024mmbench). More broadly, prior work shows that ambiguous and hard-to-solve comparative prompts can systematically induce spurious preferences in models, meaning that the evaluation prompts themselves can become a hidden source of bias when they implicitly force a choice without sufficient grounding or context (adiga-etal-2025-attention). This further motivates interventions like circular evaluations and other option-robust MCQ protocols. Other efforts focus on statistical refinement of evaluation metrics. For instance, fluid2025 apply Item Response Theory (IRT) motivated weighting to account for item difficulty and discrimination beyond simple average accuracy.
@@
Beyond these issues, multiple-choice formats are also misaligned with real-world VLM usage, where models are typically deployed in open-ended, generative settings rather than selecting from a small, predefined set of options. As a result, strong MCQ performance may overstate practical capability by rewarding option elimination or prompt-specific biases, as MCQ-based evaluations systematically misrepresent model abilities by constraining outputs and failing to probe the generative behaviors that dominate real-world LLM and VLM deployment (li2024mcqllm).
@@
Additional analyses suggest that many VLMs can perform well on certain benchmarks without meaningfully leveraging visual input, calling into question whether such evaluations truly measure visual understanding or multimodal reasoning (vblindlee-etal-2025-vlind; vblindli2024surveybenchmarksmultimodallarge; vblindlin2024revisitingrolelanguagepriors; vblindwang2024picture; vblindzhang2025debiasingmultimodallargelanguage). In contrast to approaches that seek to recover signal through post hoc statistical modeling, our method improves evaluation reliability at the source by enhancing data quality via systematic transformation and filtering of benchmark examples, building on both prior work and newly introduced techniques.
@@
Efficient & Discriminative Evaluation. Efforts to improve the efficiency of model evaluation largely draw from (1) psychometric modeling, and (2) exploiting semantic structure in evaluation data. IRT–based methods (fluid2025; polo2024tinybenchmarks) model latent capability variables in order to estimate item difficulty and discrimination. In practice, however, these approaches typically require large, dense response matrices (many models evaluated on many items) to fit parameters stably. Without this scale, estimates can become highly sensitive to hyperparameter choices.
@@
An alternative line of work leverages semantic structure. For example, vivek2024anchor employ embedding-based clustering to select representative subsets, while Scales++ (bean2025scales++) relies on qualitative, rubric-driven segmentation of tasks. These approaches face notable limitations. Clustering outcomes are tightly coupled to the choice of embedding model, a significant concern given the lack of unified multimodal embeddings, while rubric-based methods are inherently labor-intensive and subjective.
@@
More broadly, approaches that optimize solely for preserving model rankings suffer from an inherent limitation. As we show in Section 3.4, rank correlation saturates quickly and can often be achieved even by random subsets whose individual samples do not reliably discriminate between weak and strong models. Consequently, prioritizing rank stability risks overfitting to a fixed set of evaluated models without guaranteeing the quality of the underlying examples. Prior work (ghosh2025onebenchtestallsamplelevel) has also proposed aggregating heterogeneous evaluations via Plackett-Luce models, emphasizing ordinal rankings for their robustness to metric calibration issues. While this addresses the challenge of combining diverse measurements, it operates downstream of data quality, aggregating rankings over noisy or blind-solvable samples still propagates those artifacts into the final ordering.
@@ 3 The Making of DatBench
Chart ChartQA (chartqa) Standard benchmark for basic chart understanding and data extraction.
@@
ChartQA Pro (chartqapro) Challenging counterpart requiring expert-authored reasoning on complex charts.
@@
CharXiv (Descriptive & Reasoning) (charxiv) Scientific charts requiring domain-specific knowledge and terminology.
@@
InfoVQA (infovqa) Mixed-media infographics (combining dense captions with visual diagrams).
@@
Document CC-OCR (Document Parsing & KIE) (yang2024ccocrcomprehensivechallengingocr) Key Information Extraction (KIE) from structured forms and receipts.
OCR-VQA (mishraICDAR19) OCR centric Q&A on book covers.
@@
OCRBench-V2 (ocrbenchv2) Comprehensive bilingual OCR benchmark; 31 scenarios covering text recognition, localization, extraction, and reasoning.
DocVQA (mathew2021docvqadatasetvqadocument) Standard benchmark for Q&A on spatial document layouts.
@@
Scene OCR TextVQA (singh2019towards) Industry standard for recognition of text embedded in natural street scenes.
@@
MME-RW (OCR in the wild) (zhang2024mme) “In-the-wild” challenges including mobile screens and digital signage.
@@
CC-OCR (Multi-Scene OCR) (yang2024ccocrcomprehensivechallengingocr) Perspectively distorted and artistically stylized text.
@@
Math / Logic MathVista (lu2024mathvista) Broad coverage of algebraic reasoning and geometry problems.
@@
Mathverse (with & without reasoning) (mathverse) Visual math benchmark with diagram-based problems across 6 information variants; tests true visual reasoning vs. text-only deduction.
@@
MathVision (mathvision) Real math competition problems with diagrams; 16 disciplines from algebra to topology.
@@
LogicVista (xiao2024logicvistamultimodalllmlogical) Interleaved text-visual clues strictly separating logic from language priors.
@@
Spatial RealWorldQA (realworldqa2024) Physical grounding in everyday photos (depth estimation, spatial relations).
@@
MME-RW (Video Monitoring & Autonomous Driving) (zhang2024mme) Safety-critical spatial awareness (autonomous driving, remote sensing).
@@
Grounding RefCOCO (kazemzadeh-etal-2014-referitgame) General referring expressions (allows both appearance and location words).
@@
RefCOCO+ (kazemzadeh-etal-2014-referitgame) Strict appearance-based grounding (disallows spatial words like “left”).
@@
RefCOCO-g (refcocog) Long, complex syntactic descriptions (testing recursive understanding).
@@
RefCOCO-M (refcocom) Cleaned and improved version of the RefCOCO (UNC) validation split for referring expression segmentation
@@
Pixmo-Point (deitke2024molmopixmoopenweights) Precision test: requires coordinate-level localization vs. bounding boxes.
@@
Counting CountBench (paiss2023teachingclipcount) Adversarial distractors to prevent density-map estimation or guessing.
@@
TallyQA (tallyqa) Open-ended counting VQA; distinguishes simple (detection-only) vs. complex (reasoning-required) questions.
@@
Diagrams AI2D (kembhavi2016diagramworthdozenimages) Standard baseline for science and engineering schematic parsing.
@@
MME-RW (Diagram/Table) (zhang2024mme) High-resolution, complex tables found in professional reports.
@@
General MMMU-Pro (yue2025mmmuprorobustmultidisciplinemultimodal) Hardest exam-style questions for reasoning depth across disciplines.
MMBench (liu2024mmbench) Evaluates conversation fidelity and instruction following.
VQA-v2 (balanced_vqa_v2) Legacy baseline for open-ended visual questioning.
@@ 3.1 MCQ Evaluations: High Noise, Low Fidelity
Standard MCQ formats systematically overestimate model capability through two primary mechanisms: random guessing and task misalignment. First, multiple-choice questions introduce a non-trivial chance baseline ($1/N$ for $N$ options), allowing models to achieve inflated scores that add significant noise to performance metrics. This inflation is compounded when evaluating across multiple stochastic samples or models; the probability of an item appearing ”solved” by at least one of $M$ uniform random guesses grows rapidly as $1-(1-1/N)^{M}$. Second, there is a fundamental mismatch between evaluation and deployment: while most VLMs are used in generative contexts, MCQs merely test the ability to pick a candidate from a pre-defined list. This ”closed-set” evaluation fails to capture the generative reasoning required for real-world tasks and allows models to rely on superficial shortcuts or linguistic priors within the options themselves (chandak2025answer). As shown in Figure 3a, this creates a ”perceived capability” bubble in which models appear proficient in MCQ formats while failing to produce the same answers in a fully generative regime.
@@
To bridge this gap, we adopt a two-pronged strategy to ensure measured performance reflects genuine visual reasoning. Wherever viable, we transform MCQs into open-ended generative tasks by removing candidate options and requiring the model to formulate a direct response. To score these free-form outputs without the brittleness of exact-string matching, we employ an LLM-as-judge (specifically Qwen3-30B (yang2025qwen3technicalreport), a cost-effective and capable judge) to perform semantic answer matching as in chandak2025answer.
@@
For tasks where options are structurally necessary, specifically inherently discriminative questions like ”Which of the following…” where generative conversion would alter the question’s core intent, we implement Circular Evaluation (liu2024mmbench). By rotating option permutations across $N$ passes and crediting a point only if the model identifies the correct answer across all rotations, we effectively collapse the chance baseline. As shown in Figure 3b across 27 models, circular evaluation yields a steeper-than-unity slope relative to vanilla MCQ. This slope captures the persistence of the ”false floor” inherent in standard formats: while vanilla MCQs grant models a significant head start (often 20-30% accuracy) through random guessing and position bias, circular evaluation reveals that genuine reasoning capability remains near zero for these same models. The steepness of the curve illustrates that vanilla MCQ continues to significantly inflate perceived performance while true accuracy remains low ($50\%$); it is only as models achieve high-level robustness that the two metrics begin to align. By stripping away this artificial inflation, we ensure the benchmark can accurately signal the transition from zero to genuine competence, a critical signal that is otherwise obscured by the noisy MCQ baseline.

### v2

@@ 1 Introduction
Empirical evaluation is the primary mechanism through which progress in foundation models is recognized, compared, and acted upon. As machine learning has shifted from narrow, task-specific systems to general-purpose vision-language models (VLMs) with broad and compositional capabilities (Wei et al., 2022), benchmarks now play an outsized role: they define what counts as progress and directly shape how substantial computational and human resources are allocated. Evaluations are no longer a passive reporting tool but an active driver of research direction.
@@
However, modern evaluation pipelines are increasingly misaligned with the behaviors they aim to measure. As model inputs span multiple modalities and outputs become increasingly generative and stochastic, benchmarks must better disentangle genuine capabilities from superficial heuristics and inherent variance (Lu et al., 2022). While the evaluation of language-only models has received sustained methodological attention (Srivastava et al., 2023; OpenCompass, 2023), VLM evaluation remains comparatively under-examined.
@@
Recent evidence suggests that this gap has become a serious liability. Existing VLM benchmarks suffer from pervasive data quality failures, including mislabeled or ambiguous examples, questions solvable without visual input, and heavy reliance on multiple-choice formats that are not representative of downstream use-cases and are vulnerable to spurious correlations (Masry et al., 2025; Liu et al., 2024; xAI, 2024; Zhang et al., 2024b; Yang et al., 2024). These artifacts inflate reported accuracy, introduce a substantial noise floor, and reduce the signal-to-noise ratio of the evaluations. In such a regime, small improvements, often on the order of a few percent, are more plausibly explained by overfitting to benchmark idiosyncrasies than by real capability gains, rendering the research community vulnerable to hill-climbing on noise (Recht et al., 2019; Wei et al., 2023).
@@
At the same time, evaluation has become a major computational bottleneck. Running comprehensive VLM evaluation suites now consumes a nontrivial fraction of total development compute (Strubell et al., 2019). For example, during the development of OLMo3, nearly 20% of the total compute budget for the post-training phase was reportedly dedicated to evaluation alone (Lambert, 2025). This burden is amplified for VLMs by the dense visual token sequences required to represent high-resolution images and the extended reasoning traces at inference time, which can collectively exceed tens of thousands of tokens per example (Bai et al., 2025). Detailed analyses indicate that much of this cost is spent evaluating samples that are either trivial, noisy, or weakly discriminative (Schick and others, 2025; Polo et al., 2024).
@@
In this work, we argue that the design of effective evaluation should be treated as a data curation problem. Rather than repeatedly constructing new benchmarks from scratch, we propose to systematically transform and filter evaluation data to maximize faithfulness, discriminative power, and efficiency. This perspective mirrors recent successes in training data curation, in which careful data transformation and selection has produced large gains in model quality and compute efficiency (Fang et al., 2023; Joshi and Mirzasoleiman, 2023; Abbas et al., 2023; Joshi et al., 2024, 2025a, 2025b; DatologyAI et al., 2024a, b, 2025). We show that the same principles apply, with similar impact, to evaluation.
@@
First, multiple-choice formats are both unfaithful and weakly discriminative in generative settings. Converting MCQs to open-ended generation reveals large hidden capability gaps. On AI2D, for example, average accuracy drops from 77.56% to 40.53%, with the strongest MCQ model losing nearly 35 points. When generative conversion is infeasible, circular evaluation (Liu et al., 2024) collapses chance baselines and exposes similar inflation effects.
Second, many VLM benchmarks can be solved without vision. By evaluating models with the image removed, we find that over 70% of samples in VQA-v2 (Goyal et al., 2017) can be answered correctly using language priors alone. Such examples fundamentally fail to measure multimodal reasoning.
Third, low-resolution inputs and inaccurate or ambiguous annotations introduce substantial noise. Using a multi-stage filtering pipeline, we discard up to 42.07% of samples in benchmarks such as MME-RealWorld (Autonomous Driving) (Zhang et al., 2024b). In these instances, evaluation is confounded by factual labeling errors and indeterminable ground truths—where poor image quality renders the target objects unrecognizable even to a human observer—effectively precluding reliable performance assessment.
@@ 2 Related Work
Faithful Evaluation. Recent research has identified significant issues with the validity of VLM benchmarks, prompting various mitigation strategies. To address inflated performance caused by high-risk baselines in multiple-choice evaluations, several studies propose reformulating tasks into generative answer-matching settings (Chandak et al., 2025) or employing circular evaluation techniques (Liu et al., 2024). More broadly, prior work shows that ambiguous and hard-to-solve comparative prompts can systematically induce spurious preferences in models, meaning that the evaluation prompts themselves can become a hidden source of bias when they implicitly force a choice without sufficient grounding or context (Adiga et al., 2025). This further motivates interventions like circular evaluations and other option-robust MCQ protocols. Other efforts focus on statistical refinement of evaluation metrics. For instance, Schick and others (2025) apply Item Response Theory (IRT) motivated weighting to account for item difficulty and discrimination beyond simple average accuracy.
@@
Beyond these issues, multiple-choice formats are also misaligned with real-world VLM usage, where models are typically deployed in open-ended, generative settings rather than selecting from a small, predefined set of options. As a result, strong MCQ performance may overstate practical capability by rewarding option elimination or prompt-specific biases, as MCQ-based evaluations systematically misrepresent model abilities by constraining outputs and failing to probe the generative behaviors that dominate real-world LLM and VLM deployment (Li et al., 2024b).
@@
Additional analyses suggest that many VLMs can perform well on certain benchmarks without meaningfully leveraging visual input, calling into question whether such evaluations truly measure visual understanding or multimodal reasoning (Lee et al., 2025; Li et al., 2024a; Lin et al., 2024; Wang et al., 2024a; Zhang et al., 2025). In contrast to approaches that seek to recover signal through post hoc statistical modeling, our method improves evaluation reliability at the source by enhancing data quality via systematic transformation and filtering of benchmark examples, building on both prior work and newly introduced techniques.
@@
Efficient & Discriminative Evaluation. Efforts to improve the efficiency of model evaluation largely draw from (1) psychometric modeling, and (2) exploiting semantic structure in evaluation data. IRT–based methods (Schick and others, 2025; Polo et al., 2024) model latent capability variables in order to estimate item difficulty and discrimination. In practice, however, these approaches typically require large, dense response matrices (many models evaluated on many items) to fit parameters stably. Without this scale, estimates can become highly sensitive to hyperparameter choices.
@@
An alternative line of work leverages semantic structure. For example, Vivek et al. (2024) employ embedding-based clustering to select representative subsets, while Scales++ (Bean et al., 2025) relies on qualitative, rubric-driven segmentation of tasks. These approaches face notable limitations. Clustering outcomes are tightly coupled to the choice of embedding model, a significant concern given the lack of unified multimodal embeddings, while rubric-based methods are inherently labor-intensive and subjective.
@@
More broadly, approaches that optimize solely for preserving model rankings suffer from an inherent limitation. As we show in Section 3.4, rank correlation saturates quickly and can often be achieved even by random subsets whose individual samples do not reliably discriminate between weak and strong models. Consequently, prioritizing rank stability risks overfitting to a fixed set of evaluated models without guaranteeing the quality of the underlying examples. Prior work (Ghosh et al., 2025) has also proposed aggregating heterogeneous evaluations via Plackett-Luce models, emphasizing ordinal rankings for their robustness to metric calibration issues. While this addresses the challenge of combining diverse measurements, it operates downstream of data quality, aggregating rankings over noisy or blind-solvable samples still propagates those artifacts into the final ordering.
@@ 3 The Making of DatBench
Chart ChartQA (Masry et al., 2022) Standard benchmark for basic chart understanding and data extraction.
@@
ChartQA Pro (Masry et al., 2025) Challenging counterpart requiring expert-authored reasoning on complex charts.
@@
CharXiv (Descriptive & Reasoning) (Wang et al., 2024c) Scientific charts requiring domain-specific knowledge and terminology.
@@
InfoVQA (Mathew et al., 2021a) Mixed-media infographics (combining dense captions with visual diagrams).
@@
Document CC-OCR (Document Parsing & KIE) (Yang et al., 2024) Key Information Extraction (KIE) from structured forms and receipts.
OCR-VQA (Mishra et al., 2019) OCR centric Q&A on book covers.
@@
OCRBench-V2 (Fu et al., 2025) Comprehensive bilingual OCR benchmark; 31 scenarios covering text recognition, localization, extraction, and reasoning.
DocVQA (Mathew et al., 2021b) Standard benchmark for Q&A on spatial document layouts.
@@
Scene OCR TextVQA (Singh et al., 2019) Industry standard for recognition of text embedded in natural street scenes.
@@
MME-RW (OCR in the wild) (Zhang et al., 2024b) “In-the-wild” challenges including mobile screens and digital signage.
@@
CC-OCR (Multi-Scene OCR) (Yang et al., 2024) Perspectively distorted and artistically stylized text.
@@
Math / Logic MathVista (Lu et al., 2024) Broad coverage of algebraic reasoning and geometry problems.
@@
Mathverse (with & without reasoning) (Zhang et al., 2024a) Visual math benchmark with diagram-based problems across 6 information variants; tests true visual reasoning vs. text-only deduction.
@@
MathVision (Wang et al., 2024b) Real math competition problems with diagrams; 16 disciplines from algebra to topology.
@@
LogicVista (Xiao et al., 2024) Interleaved text-visual clues strictly separating logic from language priors.
@@
Spatial RealWorldQA (xAI, 2024) Physical grounding in everyday photos (depth estimation, spatial relations).
@@
MME-RW (Video Monitoring & Autonomous Driving) (Zhang et al., 2024b) Safety-critical spatial awareness (autonomous driving, remote sensing).
@@
Grounding RefCOCO (Kazemzadeh et al., 2014) General referring expressions (allows both appearance and location words).
@@
RefCOCO+ (Kazemzadeh et al., 2014) Strict appearance-based grounding (disallows spatial words like “left”).
@@
RefCOCO-g (Mao et al., 2016) Long, complex syntactic descriptions (testing recursive understanding).
@@
RefCOCO-M (moondream, 2025) Cleaned and improved version of the RefCOCO (UNC) validation split for referring expression segmentation
@@
Pixmo-Point (Deitke et al., 2024) Precision test: requires coordinate-level localization vs. bounding boxes.
@@
Counting CountBench (Paiss et al., 2023) Adversarial distractors to prevent density-map estimation or guessing.
@@
TallyQA (Acharya et al., 2019) Open-ended counting VQA; distinguishes simple (detection-only) vs. complex (reasoning-required) questions.
@@
Diagrams AI2D (Kembhavi et al., 2016) Standard baseline for science and engineering schematic parsing.
@@
MME-RW (Diagram/Table) (Zhang et al., 2024b) High-resolution, complex tables found in professional reports.
@@
General MMMU-Pro (Yue et al., 2025) Hardest exam-style questions for reasoning depth across disciplines.
MMBench (Liu et al., 2024) Evaluates conversation fidelity and instruction following.
VQA-v2 (Goyal et al., 2017) Legacy baseline for open-ended visual questioning.
@@ 3.1 MCQ Evaluations: High Noise, Low Fidelity
Standard MCQ formats systematically overestimate model capability through two primary mechanisms: random guessing and task misalignment. First, multiple-choice questions introduce a non-trivial chance baseline ($1/N$ for $N$ options), allowing models to achieve inflated scores that add significant noise to performance metrics. This inflation is compounded when evaluating across multiple stochastic samples or models; the probability of an item appearing ”solved” by at least one of $M$ uniform random guesses grows rapidly as $1-(1-1/N)^{M}$. Second, there is a fundamental mismatch between evaluation and deployment: while most VLMs are used in generative contexts, MCQs merely test the ability to pick a candidate from a pre-defined list. This ”closed-set” evaluation fails to capture the generative reasoning required for real-world tasks and allows models to rely on superficial shortcuts or linguistic priors within the options themselves (Chandak et al., 2025). As shown in Figure 3a, this creates a ”perceived capability” bubble in which models appear proficient in MCQ formats while failing to produce the same answers in a fully generative regime.
@@
To bridge this gap, we adopt a two-pronged strategy to ensure measured performance reflects genuine visual reasoning. Wherever viable, we transform MCQs into open-ended generative tasks by removing candidate options and requiring the model to formulate a direct response. To score these free-form outputs without the brittleness of exact-string matching, we employ an LLM-as-judge (specifically Qwen3-30B (Yang et al., 2025), a cost-effective and capable judge) to perform semantic answer matching as in Chandak et al. (2025).
@@
For tasks where options are structurally necessary, specifically inherently discriminative questions like ”Which of the following…” where generative conversion would alter the question’s core intent, we implement Circular Evaluation (Liu et al., 2024). By rotating option permutations across $N$ passes and crediting a point only if the model identifies the correct answer across all rotations, we effectively collapse the chance baseline. As shown in Figure 3b across 27 models, circular evaluation yields a steeper-than-unity slope relative to vanilla MCQ. This slope captures the persistence of the ”false floor” inherent in standard formats: while vanilla MCQs grant models a significant head start (often 20-30% accuracy) through random guessing and position bias, circular evaluation reveals that genuine reasoning capability remains near zero for these same models. The steepness of the curve illustrates that vanilla MCQ continues to significantly inflate perceived performance while true accuracy remains low ($50\%$); it is only as models achieve high-level robustness that the two metrics begin to align. By stripping away this artificial inflation, we ensure the benchmark can accurately signal the transition from zero to genuine competence, a critical signal that is otherwise obscured by the noisy MCQ baseline.